Tohban Report 2008-10-29
From RHESSI Wiki
|Start Date:||22 Oct 2008|
|End Date:||29 Oct 2008|
|Next Tohban:||Albert Shih|
|List all reports|
Very quiet. GOES only once reached A level (briefly).
SSR is emptying with each pass, thanks to more (~10) passes per day
Cold plate 2 is still rising, now at 97.7 K.
Data Gaps in VC3-MONITOR RATES:
- 2008-10-27T02:30:00.000 -- 2008-10-27T02:40:00.000 600.00000
On Thursday, the spacecraft was spun up from 14.5 to 15 nominal, currently at 14.9. Front segment data are not being taken during eclipse.
Some comments from Dave Smith: Detectors 6 and 8 are still worsening. Options: Continue to lower HV, which we should do within the next few weeks if we choose this option; or, 'go for broke' by raising the HV instead.
Sam Krucker gave a science talk entitled "RHESSI and STEREO observations of a partially disk-occulted flare".